OpenStudy (anonymous):

is it necessary to write equivalent fractions using the lcd if you are subtracting 13/4 from 22/3? please explain

OpenStudy (anonymous):

Not necessary, though if you're doing it by hand, it might be easier. If you have a calculator, just go ahead and type in 22/3 - 13/4, if you're doing it by hand: 22/3 = (22*4) / (3*4) and 13/4 = (13*3)/(4*3) So then you can do: (22*4)/(3*4) - (13*3)/(4*3) = 88/12 - 39/12 = 49/12